Find the mass of the object, which is in the form of a sphere of radius √5cm, centred at the origin. The density at any point is given to be constant 2.


1
Expert's answer
2021-04-14T14:58:17-0400

Mass of the object "m" could be found by multiplying volume "V" by density "\rho".

"V=\frac{4}{3}\pi r^3 = \frac{4}{3}\pi 5^{\frac{3}{2}}" - volume of sphere

"m = V\rho=\frac{4}{3}\pi 5^{\frac{3}{2}}*2 = \frac{8}{3}\pi 5^{\frac{3}{2}}"